OpenStudy (anonymous):

(9sqrtx^7)(9sqrtx^5)

OpenStudy (anonymous):

\[(9\sqrt{x ^{7}})(9\sqrt{x ^{5}})\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

its one top of the other. i got mixed up haha

OpenStudy (nincompoop):

\[9\sqrt[7]{x} \times 9\sqrt[5]{x}\]

OpenStudy (nincompoop):

like that?

OpenStudy (anonymous):

the second part is in the bottom

OpenStudy (nincompoop):

\[\frac{ 9\sqrt{x^7} }{ 9 \sqrt{x^5} }\] or like this?

OpenStudy (anonymous):

like thast!

OpenStudy (anonymous):

that*

OpenStudy (nincompoop):

how do you think you will solve this? and do you already know the rules of exponents?

OpenStudy (anonymous):

i have no idea. thats why i need help /:

OpenStudy (nincompoop):

awww do you have a book or something?

OpenStudy (anonymous):

noope

OpenStudy (nincompoop):

can you maybe read the content of this link first and then figure out how to answer your question? http://www.mathsisfun.com/algebra/exponent-laws.html and if you are stuck, show us what you have as a solution then we will guide you. If I taught you how to tackle this problem now, it wouldn't really make any sense because you haven't got the fundamental knowledge required yet.

OpenStudy (anonymous):

okay thank you so much!